OpenStudy (abymartinez15):

can someone please help me A company makes and sells charm bracelets. The cost of producing x bracelets is represented by the function C(x) = 180 + 8x. The revenue earned from selling x bracelets is represented by the function R(x) = 20x. Write and simplify a function P that represents the profit made from selling x bracelets. How many bracelets must the company sell to break even?

OpenStudy (rayaz):

Profit = Revenue - Cost so subtract C(x) from R(x)

OpenStudy (abymartinez15):

okay

OpenStudy (abymartinez15):

so C(x) would be 180?

OpenStudy (rayaz):

let the function P(x) represent the profit earned. P(x) = R(x) - C(x) P(x) = 20x - (180 + 8x) P(x) = 20x -180 -8x P(x) = 20x - 8x -180 P(x) = 12x - 180

OpenStudy (abymartinez15):

so i subtract 12-180?

OpenStudy (rayaz):

No. You have to find the value of x when the company breaks even (i.e they do not suffer losses nor do they make a profit.)

OpenStudy (abymartinez15):

oh okay

OpenStudy (abymartinez15):

-168?

OpenStudy (rayaz):

No. Set P(x) = 0 (This means that they do not make a profit but do not suffer a loss either.) 0 = 12x -180 12x = 180 x = 15 So they need to manufacture 15 bracelets to break even.

OpenStudy (abymartinez15):

okay.. so it will be 12(15)=180 right?

OpenStudy (rayaz):

Yes

OpenStudy (abymartinez15):

thank you